Projection Admin 5/9
To do this evening/ was completed: The rain has been persistent, so we'll be late on changing the marquee and posters. Hopefully it'll subside by tomorrow morning.
This morning Jack was having issues with the computer in 4/6 as all external drives were readable but not writeable. All of these drives were formatted as NTFS and I've been trying to figure out if the OS update has anything to do with compatibility for reading/writing, because they should kind of work in that state on a mac (but some evidence online suggests otherwise and it says Windows NTFS so, you know, that's not ideal) and they want you to use APFS for Ventura (which is what the computer is on now) but exFAT is what we want to use the drives on both macs and PCs. We didn't need anything off the George Ivan drive so I reformatted it for exFAT which made it readable/writeable. The other two external drives, preshows/trailers and VHS Vault, are still NTFS. I copied over the Alice Jr. DCP to George Ivan in a new folder entitled Current_SE_Features (we have this ingested but it won't screen for a while and if we need to clear content off the servers in the meantime, I'd like it to be on a drive somewhere), along with a Current Trailers folder (nothing in there right now. we don't really need anything in there), all the PSA/Policies have been copied over, and now I'm just copying over the Generic Preshows. Once those are done, we can copy over the SE/Brunch preshows. After all that's sorted, my plan is to reformat the preshows/trailers drive and put everything back. This is, basically, the only way we can format it properly and still keep all that stuff downloaded. Not tough though! Just transferring. ALSO in the meantime I just downloaded our new content from Kat to the desktop and ingesting it through the SMB. Easy enough. Please continue working on the new trailer list tonight. Thank you!
The Nicholas Winding Refin Drive I absolutely cannot figure out still, but that's for another time.
Presentation Issues:
Preshow Clips/Playlist: Tech'd Hairspray cue sheets ever so slightly updated along with the playlist.
Presentation Issues:
Preshow Clips/Playlist: Tech'd Hairspray cue sheets ever so slightly updated along with the playlist.
Trailer Updates:
Special Events/ AV set up: I took a look at the Make MKV file vs the Handbrake file and the Make MKV looked way better so I've kept that one ( only thing is the dimensions), but stopped the DCP. I reached out to Chenelle to see if we could get a BluRay. Apparently that's what she thought they bought, but they did not, they bought the DVD.
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content): Jack ingested Paris is Burning/ The Howling
Downloaded and ingested the following content (to the desktop in 4/6) from Kat:
TLR_BCB_2025 (On the drive)
TLR_NPP_FORECAST_0516_A
TLR_NPP_FORECAST_0516_B
PS_SE_ReconsiderThis_V2_5min_0525
TLR_NPP_Matrix Revolutions
Outgoing/Shipping (content):
Print/video Inspection: Tech'd all of the NoBudge shorts with Desmond today.
Tech'd Bitter Tears with Cristina. For some odd reason, P1 optical sound wasn't initially working. Everything looked fine; connections secure, alignment/sound head, processor, etc. I blew some canned air, threaded up another reel, changed over twice and it worked. Weird! Something was off and struggling but I'm not sure what just yet. I kept the reels running and changed over a few more times just to make sure.
Jack ran Barbershop in T4.
Asset tracking doc updated/checked?:
Maintenance/cleaning:
Cue sheets updated/checked (every shift): Yes
Marquee updated?:
Posters updated?:
Cleaning lights on for the night?:
Projection cart by entrance?:
Additional Notes:
Comments
Post a Comment